OpenStudy (anonymous):

Solve the equation by expressing each side as a power of the same base. e^x+9=1/e^7 a. 16 b. -2 c. 2 d. -16

OpenStudy (zehanz):

I think you mean:\[e^{x+9}=\frac{1}{e^7}\] Now remember this rule:\[\frac{1}{a^b}=a^{-b}\]

OpenStudy (anonymous):

Thats exactly the problem but I was never taught a rule?

OpenStudy (anonymous):

Its just solving logarithmic equation using same base

OpenStudy (zehanz):

You're taught now... so:\[e^{x+9}=e^{-7} \Leftrightarrow x+9 = -7\]Solve it to find x.